The single payment compound amount factor is used to find the value of sum and is written as (F/P, i, n) O present...recurring O present.future O future...recurring O future...present
Answer #1

Solution -

Option D. Future...Present is Correct Option

Reason -

in Single Payment Compund Amount Factor

When Find F

Given P

F= P( 1 + i) ^n = P ( F/P,i,n)

So here

Option D. Future...Present is Correct Option

NOTE - DO UP VOTE HAVE A NICE DAY